From a group of 12 people I am to form 4 teams of 3 players each. In how many ways can this be done. (This is closely related to the previous problem, but note that the teams are not ordered.
I tried doing (12x11x10x9)/(3x2x1)
and
12!/4!3!
but both are wrong

Answer by stanbon(75887)

From a group of 12 people I am to form 4 teams of 3 players each. In how many ways can this be done.
---
# of ways to form the 1st team of 3: 12C3 = (12*11*10)/(1*2*3) = 220 ways
---
# of ways to form the 2nd team: 9C3 = (9*8*7)/(1*2*3) = 84 ways
---
# of ways to form the 3rd team: 6C3 = (6*5*4)/(1*2*3) = 20 ways
---
# of ways to form the 4th team: 3C3 = 1 way
------------------------------------------------
Total # of ways: 220*84*20*1 = 369,600
===========
Cheers,
Stan H.